If I purchase a home for $150,000. I do a fixed 30-year mortgage at 6%. My total fees and percent down on the home equaled $12,500. What would be my annual payment?

a. $12,345.67

b. $9,376.24

c. $6,456.72

d. $9,712.72
Answer #1

Annual Payments = P*r*(1 + r) (1+r) - 1

where n is the no of years

P =  Principle - Down payment

r is the rate of interest per annum

Annual Payments = 137,500 * 0.06*(1 +0.06) 30 (1 + 0.0630-1

= 47383.98022/ 4.7434912

= $ 9989.225

The correct answer is $ 9989.225 which is not in Option. I have checked all the calculations 3 times. There is no error.

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.